Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

random sum equals 9

Expert replies
by EricKryk » Fri Jan 24, 2014 9:30 am
A = {2, 3, 4, 5}
B = {4, 5, 6, 7, 8}

Two integers will be randomly selected from the sets above, one integer from set A and one integer from set B. What is the probability that the sum of the two integers equal 9?

A) 0.15
B) 0.20
C) 0.25
D) 0.30
E) 0.33
Join the discussion
Source: — Problem Solving |

by Brent@GMATPrepNow » Fri Jan 24, 2014 9:37 am
EricKryk wrote:A = {2, 3, 4, 5}
B = {4, 5, 6, 7, 8}

Two integers will be randomly selected from the sets above, one integer from set A and one integer from set B. What is the probability that the sum of the two integers equal 9?

A) 0.15
B) 0.20
C) 0.25
D) 0.30
E) 0.33
Here's a quick solution.
First randomly select a number from set A.
At this point, you can see that, of the 5 numbers in set B, only 1 of them will combine with your number from set A to create a sum of 9.

So, the probability of getting a sum of must = [spoiler]1/5 = 0.2 = B[/spoiler]

Cheers,
Brent
Last edited by Brent@GMATPrepNow on Mon May 18, 2015 6:04 am, edited 2 times in total.
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Patrick_GMATFix » Fri Jan 24, 2014 9:43 am
Find out the number of pairs that add up to 9 and put that over all possible pairs to find the requested probability. The solution below is taken from the GMATFix App.

Image

-Patrick
  • Ask me about tutoring.
Join the discussion

by [email protected] » Fri Jan 24, 2014 6:55 pm
Hi EricKryk,

Probability questions are based on the probability formula:

(Number of ways that you "want") / (Total number of ways possible)

Since it's usually easier to calculate the total number of possibilities, I'll do that first. There are 4 options for set A and 5 options for set B; since we're choosing one option from each, the total possibilities = 4 x 5 = 20

Now, to figure out the number of duos that sum to 9:

2 and 7
3 and 6
4 and 5
5 and 4

4 options that give us what we "want"

4/20 = 1/5 = 20% = .2 = B

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by sanju09 » Fri Jan 24, 2014 10:39 pm
EricKryk wrote:A = {2, 3, 4, 5}
B = {4, 5, 6, 7, 8}

Two integers will be randomly selected from the sets above, one integer from set A and one integer from set B. What is the probability that the sum of the two integers equal 9?

A) 0.15
B) 0.20
C) 0.25
D) 0.30
E) 0.33
For any integer selected from the 4 in set A, we've 5 ways to club it with an integer from set B. Hence, the total number of ways to select two integers this way are (4)(5) = 20. The favorable ways are 2 + 7, 3 + 6, 4 + 5, and 5 + 4, 4 in number.

Required probability = [spoiler]4/20 or 0.20.

Pick B
[/spoiler]
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion

by Jeff@TargetTestPrep » Mon May 18, 2015 5:55 am
EricKryk wrote:A = {2, 3, 4, 5}
B = {4, 5, 6, 7, 8}

Two integers will be randomly selected from the sets above, one integer from set A and one integer from set B. What is the probability that the sum of the two integers equal 9?

A) 0.15
B) 0.20
C) 0.25
D) 0.30
E) 0.33
Solution:

To determine the probability that the sum of the two integers will equal 9, we must first recognize that probability = (favorable outcomes)/(total outcomes).

Let’s first determine the total number of outcomes. We have 4 numbers in set A, and 5 in set B, and since we are selecting 1 number from each set, the total number of outcomes is 4 x 5 = 20.

For our favorable outcomes, we need to determine the number of ways we can get a number from set A and a number from set B to sum to 9. We are selecting from the following two sets:

A = {2, 3, 4, 5}

B = {4, 5, 6, 7, 8}

We will denote the first number as from set A and the second from set B. Here are the pairings that yield a sum of 9:

2,7
3,6
4,5
5,4

We see that there are 4 favorable outcomes. Thus, our probability is 4/20 = 0.25, Answer C.

Jeffrey Miller
Head of GMAT Instruction
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews
Join the discussion